# Solving the Gross-Pitaevskii Equation with Quantic Tensor Trains: Ground States and Nonlinear Dynamics

### v2

@@ Abstract
We develop a tensor network framework based on the quantic tensor train (QTT) format to efficiently solve the Gross-Pitaevskii equation (GPE), which governs Bose-Einstein condensates under mean-field theory. By adapting time-dependent variational principle (TDVP) and gradient descent methods, we accurately handle the GPE’s nonlinearities within the QTT structure. Our approach enables high-resolution simulations with drastically reduced computational cost. We benchmark ground states and dynamics of BECs–including vortex lattice formation and breathing modes–demonstrating superior performance over conventional grid-based methods and stable long-time evolution due to saturating bond dimensions. This establishes QTT as a powerful tool for nonlinear quantum simulations.
@@ 1 Introduction
Efficient numerical representations of high-dimensional data are essential in both quantum many-body physics and applied mathematics [1, 2]. Tensor network methods provide a systematic approach to compressing such data while preserving essential correlations. Matrix product states (MPS), in particular, have been widely used to approximate low-entanglement quantum states, enabling accurate simulations of low-dimensional quantum systems [3, 4]. In applied mathematics, an analogous approach is the tensor train (TT) decomposition, which shares the same mathematical structure as MPS and is widely used for efficiently representing high-dimensional data [5, 6, 7].
@@
A key extension of tensor train decomposition is the quantic tensor train (QTT), a quantum-inspired numerical method that efficiently discretizes continuous functions on an exponential grid. A QTT maps high-dimensional tensors into a logarithmic-size tensor train, which significantly reduces storage requirements and computational complexity from exponential to polynomial [8, 9]. Unlike conventional grid-based discretization techniques, QTT exploits a hierarchical structure similar to quantum wavefunction representations, where different ”qubits” correspond to different length scales. Consequently, the tensor ranks for smooth functions are typically small. The ability to encode large-scale problems makes QTT powerful for various applications, including Fourier transforms [10, 11], quantum field theory [12, 13, 14], partial differential equations [15, 16], quantum chemistry [17], and data compression [18, 19].
@@
Since QTT can efficiently represent continuous-space functions and operators, it is particularly useful for solving quantum systems involving single particles or systems under the mean-field approximations [20, 21]. Given that MPS and QTT essentially share the same tensor network structure but differ in their applications, numerical algorithms developed for many-body wavefunctions, such as the density matrix renormalization group (DMRG) [3, 4], can be applied naturally to QTT wavefunctions. Similarly, time evolutions of QTT wavefunctions can be effectively performed using the time-dependent variational principle (TDVP) [22, 23].
@@
In this work, we are particularly interested in the applications of solving the Gross-Pitaevskii equation (GPE)[24], a non-linear Schrödinger equation describing behaviors of Bose-Einstein condensates (BECs), where a large fraction of particles condense onto a single quantum state [25], under the mean-field approximation. BECs have been experimentally realised in the highly tunable ultra-cold dilute gas systems in a variety of settings with the state-of-the-art experimental technologies, e.g. [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. This allows researchers to study the dynamics of macroscopic wavefunction, superfluidity, collective excitations, topological excitations, superfluid turbulences and etc. [28], which have been successfully captured by the GPE. BEC systems are closely related superfluidity, and the emergence of quantum vortices is one of the hallmarks of that. Vortices are the angular momentum carriers in superfluid systems and can be generated by, for example, rotation [29], external stirring [32] or atom-light coupling [35]. For large-scale systems accommodating multiple vortices, accurate numerical simulations require sufficiently fine spatial discretisation, presenting significant computational challenges. Similar difficulties arise in large-scale simulations of nonlinear Schrödinger equations in cosmological [38, 39, 40, 41] and astrophysical systems [42, 43, 44]. The QTT framework effectively addresses these issues by enabling an exponential increase in discretization points.
@@
Due to the intrinsic nonlinearity of the GPE, standard MPS methods, such as DMRG and TDVP, cannot be applied directly without modifications. We extend existing MPS algorithms to incorporate the nonlinearities of the GPE, allowing us to efficiently compute its ground states and dynamical evolution. To find the ground states, we consider two numerical approaches: imaginary-time evolution (ITE), also called as the imaginary-time propagation method [45], and variational methods. We benchmark our methods by considering a cylindrically harmonically trapped BEC in two dimensions with and without rotation for ground-state solutions and the dynamics of its breathing-mode (mono-pole) excitation. Rotating a BEC results in the formation of quantum vortices, which organise themselves into a triangular lattice format, similar to the Abrikosov lattice in superconductors [46]. On the other hand, the breathing-mode excitation is a BEC width oscillation with a trapping-frequency-dependent frequency for a sufficiently large BEC [47, 48, 49], and we generate it by a sudden quench in the interaction strength.
@@
We note that similar ideas have been explored independently in Ref. [50, 51, 52], which appeared concurrently with our work.
@@ 2 Gross-Pitaevskii equation
GPE is a nonlinear Schrödinger equation describing the weakly interacting BEC under mean-field approximation. In this work, we consider BEC in a rotating harmonic trap [53]. The Hamiltonian is given by
@@
$\hat{H}_{\mathrm{int}}=g|\psi(\mathbf{r},t)|^{2}$ (3)
where $\hat{H}^{1}$ is the one-body Hamiltonian including the kinetic energy, the harmonic trapping potential $\frac{1}{2}m\omega^{2}r^{2}$, and the rotational term $\Omega\hat{L}_{z}$. $\hat{H}_{\mathrm{int}}$ is the interaction with strength $g=4\pi\hbar^{2}a_{s}N/m$, where $a_{s}$ is the $s$-wave scattering length, $N$ is the total particle number and $m$ is the particle mass. Throughout this work, we consider the dimensionless Hamiltonian by adopting the characteristic scales of quantum harmonic oscillator units, $\sqrt{\hbar/m\omega}$, $1/\omega$ and $\hbar\omega$ for the length, time and energy respectively (This is equivalent to set $\hbar=1$, $m=1$, and $\omega=1$). The condensate wavefunction is normalized to one. The evolution of the wavefunction is governed by the time-dependent GPE
$i\hbar\frac{\partial\psi(\mathbf{r},t)}{\partial t}=\hat{H}\psi(\mathbf{r},t)$ (4)
@@
$E^{1}[\psi]=\int d^{3}\mathbf{r}\,\psi^{*}\hat{H}^{1}\psi(\mathbf{r})\;;$ (7)
$E_{\mathrm{int}}[\psi]=\frac{g}{2}N\int d^{3}\mathbf{r}|\psi(\mathbf{r})|^{4}.$ (8)
@@ 3 Methods
In this section, we introduce two methods for computing the ground state of the GPE, including the ITE and a gradient descent method. For real- or imaginary-time evolution, TDVP is a powerful method with the MPS/QTT framework, and GD is effective for variationally optimizing the ground state. Here, we explain the necessary modifications to both methods to handle the nonlinear interaction term in the Hamiltonian.
@@ 3.1 Imaginary-time evolution: TDVP
ITE is a common strategy for computing the ground state of the GPE [45] and can be applied to the QTT wavefunction with minor modifications. Given the wavefunction (Hamiltonian) as a QTT (QTTO), the time evolution of a small time step $\delta\tau$ can be performed by using TDVP algorithm [22, 23]. Although the Hamiltonian in QTTO representation is typically “long-range” in the sense that it couples qubits over large distances, TDVP operates directly on the QTTO Hamiltonian, which makes it particularly well suited for QTT time evolutions. The Hamiltonian of the GPE contains the one-body term $\hat{H}^{1}$ and the interaction term $\hat{H}_{\mathrm{int}}$. The interaction term $gN|\psi(\mathbf{r})|^{2}$ depends on the wavefunction $\psi(\mathbf{r})$, and therefore needs to be updated at each step using the evolved wavefunction from the last step. The time step $\delta\tau$ is set to be $0.5\delta x^{2}$ to satisfy the Courant-Friedrichs-Lewy condition and ensure the stability of the convergence.
An essential step in the TDVP algorithm is to compute the effective Hamiltonian, where the required tensor contractions are shown in Fig. 1(a). The computational cost is proportional to $dD^{3}D_{W}+d^{2}D^{2}D_{W}^{2}$, where $d=2$ is the dimension for each qubit, and $D$ and $D_{W}$ are the bond dimensions of the wavefunction and the Hamiltonian. We keep $\hat{H}^{1}$ and $\hat{H}_{\mathrm{int}}$ as two QTTOs and compute their effective Hamiltonians separately. For $\hat{H}^{1}$, the bond dimension is fixed and small, and the corresponding computational cost is relatively low. However, the QTTO for the interaction has a bond dimension $D^{2}$ for an exact calculation, where $D$ is the bond dimension of the wavefunction. Since $D$ controls the accuracy of the wavefunction and can be large in a challenging system, we need to compress $\psi^{2}(\mathbf{r})$ from a bond dimension $D^{2}$ to a smaller bond dimension $\chi$ to achieve efficient computations (see Fig. 1(b)). The computational cost then depends on the efficiency of this compression. In this work, we perform the compression using the variational method by maximizing the overlap $\langle\psi^{2}|\psi^{2}_{\chi}\rangle$ [2], where $\psi^{2}_{\chi}(\mathbf{r})$ is the compressed $\psi^{2}(\mathbf{r})$ with the bond dimension $\chi$. In practice, $\psi^{2}(\mathbf{r})$ typically has a structure similar to that of $\psi(\mathbf{r})$, and therefore $\chi$ is typically on the same order as $D$. This yields an overall computational complexity of roughly $O(D^{4})$.
@@ 3.2 Variational method: gradient descent
Another powerful approach to the ground state is to use the variational method to minimize the energy $E[\psi]$. Although DMRG is not applicable due to the nonlinearity of the Hamiltonian, gradient descent remains a viable optimization method. Gradient descent has been used to optimize an MPS when the tensors are isometry tensors [54, 55, 56, 57]. In this work, we employ gradient descent within the MPS/QTT framework. The algorithm proceeds as follows. The QTT wavefunction is represented in canonical form, with one tensor at the orthogonality center–initially at the first site–and all tensors to the left (right) being left (right) orthogonal. The tensor at the orthogonality center is updated based on its energy gradient until it converges to a minimum. Once optimized, the orthogonality center is moved to the next site, and the process is repeated. This sweep continues back and forth until the global energy minimum is reached.
@@
Figure 2: Errors in the gradient of the interaction energy (black curve) and in the slope along the descent direction (red curve) as functions of the number of tensors updated in $\psi_{\chi}^{2}$ after updating a tensor in $\psi$.

### v3

@@ Abstract
We develop a tensor network framework based on the quantic tensor train (QTT) format to efficiently solve the Gross-Pitaevskii equation (GPE), which governs Bose-Einstein condensates under mean-field theory. For sufficiently regular functions, the QTT approach enables exponentially fine grids with computational cost that grows only linearly. By adapting the time-dependent variational principle (TDVP) and gradient descent methods, we accurately handle the GPE’s nonlinearities within the QTT structure. Our approach enables high-resolution simulations with drastically reduced computational cost on highly refined grids. We benchmark the ground states and dynamics of BECs, including vortex lattice formation and breathing modes, demonstrating superior performance over regular-grid-based methods and stable long-time evolution due to saturating bond dimensions. This establishes QTT as a powerful tool for nonlinear quantum simulations.
@@ 1 Introduction
Efficient numerical representations of high-dimensional data are essential in both quantum many-body physics and applied mathematics [McCulloch_2007, SCHOLLWOCK201196]. Tensor network methods provide a systematic approach to compressing such data while preserving essential correlations. Matrix product states (MPS), in particular, have been widely used to approximate low-entanglement quantum states, enabling accurate simulations of low-dimensional quantum systems [PhysRevB.48.10345, PhysRevLett.69.2863]. In applied mathematics, an analogous approach is the tensor train (TT) decomposition, which shares the same mathematical structure as MPS and is widely used for efficiently representing high-dimensional data [doi:10.1137/090752286, Oseledets2013, Khoromskij+2018].
@@
A key extension of tensor train decomposition is the quantic tensor train (QTT), a quantum-inspired numerical method that efficiently discretizes continuous functions on an exponential grid. A QTT maps high-dimensional tensors into a logarithmic-size tensor train, which significantly reduces storage requirements and computational complexity from exponential to polynomial [Khoromskij2011, Oseledets2009]. Unlike conventional grid-based discretization techniques [^1] In our paper, the ”conventional” specifically refers to the method with regular finite-difference discretization, which, at least conceptually, saves the data of all the grid points in memory., QTT exploits a hierarchical structure similar to quantum wavefunction representations, where different ”qubits” correspond to different length scales. Consequently, the tensor ranks for smooth functions are typically small. The ability to encode large-scale problems makes QTT powerful for various applications, including Fourier transforms [Dolgov2012, PRXQuantum.4.040318], quantum field theory [PhysRevX.13.021015, PhysRevX.12.041018, 10.21468/SciPostPhys.18.1.007], partial differential equations [khoromskij2014tensor, garcia2021quantum], quantum chemistry [jolly2024tensorizedorbitalscomputationalchemistry], and data compression [PhysRevResearch.4.043007, PhysRevLett.132.056501].
@@
Since QTT can efficiently represent continuous-space functions and operators, it is particularly useful for solving quantum systems involving single particles or systems under the mean-field approximations [garcia2023global, li2024matrixproductstatesquantization]. Given that MPS and QTT essentially share the same tensor network structure but differ in their applications, numerical algorithms developed for many-body wavefunctions, such as the density matrix renormalization group (DMRG) [PhysRevB.48.10345, PhysRevLett.69.2863], can be applied naturally to QTT wavefunctions. Similarly, time evolutions of QTT wavefunctions can be effectively performed using the time-dependent variational principle (TDVP) [PhysRevLett.107.070601, PhysRevB.94.165116].
@@
In this work, we are particularly interested in the applications of solving the Gross-Pitaevskii equation (GPE) [rogel2013gross], a non-linear Schrödinger equation describing behaviors of Bose-Einstein condensates (BECs), where a large fraction of particles condense onto a single quantum state [einstein1924quantum], under the mean-field approximation. BECs have been experimentally realised in the highly tunable ultra-cold dilute gas systems in a variety of settings with the state-of-the-art experimental technologies, e.g. [Lin2003kih, doi:10.1126/science.269.5221.198, RevModPhys.71.463, abo2001observation, Hadzibabic:2006jjp, Chin2010RMP, Madison2000PhysRevLett.84.806, Gerbier:2010zz, Donadello2014, PhysRevLett.83.2498, Karailiev:2024sxs, Geng:2025uup]. This allows researchers to study the dynamics of macroscopic wavefunction, collective quasi-particle excitations , topological excitations, quantum turbulences, etc. [RevModPhys.71.463, Barenghi2023quantum, Proukakis2025], which have been successfully captured by the GPE. BEC systems are closely related superfluidity, and the emergence of quantum vortices is one of the hallmarks of that. Vortices are the angular momentum carriers in superfluid systems and can be generated by, for example, rotation [abo2001observation], external stirring [Madison2000PhysRevLett.84.806] or atom-light coupling [PhysRevLett.83.2498]. In addition, the high tunability, precise controllability, and advanced experimental techniques make BEC systems ideal candidates for quantum simulation [RevModPhys.86.153, Proukakis2025] and emerging atomtronic applications [Amico2021roadmap, Proukakis2025].
Typically, there are two types of problems in solving the GPE: the nonlinear eigensolutions and the time propagation of it (see Ref. [Minguzzi2004] for a comprehensive review). A recent review for the former can be found in Ref. [Henning2025], and it is often referred to as the lowest-energy solution of a GPE [bao2003numerical]. A common method is known as the imaginary-time evolution/propagation (ITE) method [Modugn2003EPJD...22..235M, Minguzzi2004], evolving the wavefunction toward the lowest energy solution as the evolution of GPE in the imaginary time is sufficiently long, but it often converges slowly and needs attention for the numerical accuracy when local minimum and energy degeneracy appear. Hence, the gradient-flow [Winiecki2001, bao2003numerical, Bland2015, Danaila2017, Chen2023, GAIDAMOUR2021108007] and Newton [Xu2013, wu2017regularized] methods have been adapted as more efficient and accurate approaches and have potential to study excited BEC states, showing fascinating physical features, such as solitonic solutions [Bland2015], enginered quantum vortices and high-orbit BEC in optical lattices [Xu2013, You2016]. The spatial derivatives in the GPE can be coped with by the finite difference and pseudo-spectral methods [GAIDAMOUR2021108007, antoine2014gpelab]. The common bases in the pseudo-spectral methods are the plane-wave and Hermite-Gaussian basese [Blakie2008, Rooney2024, Keepfer2022], depending on setups and system geometries. For the cases of finite difference and Fourier pseudo-spectral method, a homogeneous meshgrid is generally required, and hence one needs sufficient grid resolution to resolve the characteristic length scale to avoid numerical errors. Adaptive meshgrid method has also been implemented for a GPE-like nonlinear equation [Schive:2014dra].
The time-evolution problem is commonly associated with the ITE method and studies on fundamental excitations of BECs [RevModPhys.71.463], nonlinear dynamics, such as mode-mixings [Morgan1998], superfluid turbulences and non-equilibrium dynamics [Schmied2019]. A common challenge is the requirement for long-time propagation in a large grid, e.g. [Schive:2014dra, Stagg2017, Schmied2019, May2021, Liu2024uvo]. The backward/forward Euler [BaoCNFD, antoine2014gpelab], time-splitting [bao2003numerical, Schive:2014dra, Symes2016, Schmied2019], Runge-Kutta and Crank-Nicolson [Ruprecht1995, Symes2016, kumar2019c, muruganandam2009fortran, antoine2014gpelab] methods have been adapted to solve the GPE.
A variety of GPE solvers exists in the literature, e.g., BEC2HPC [GAIDAMOUR2021108007], GPELab [antoine2014gpelab, antoine2015gpelab], GPUE [PhysRevA.94.053603], BEC-GP-ROT-OMP [kumar2019c], in addition to the works addressed previously and their associated references. However, for large-scale systems accommodating multiple vortices, accurate numerical simulations require sufficiently fine spatial discretisation, presenting significant computational challenges. Similar difficulties arise in large-scale simulations of nonlinear Schrödinger equations in cosmological [Schive:2014dra, PhysRevD.84.043531, Huang:2022ffc, Chan:2025hhg] and astrophysical systems [Graber2016imq, Wood:2022smd, Liu2024uvo], where physics at vastly different length scales are importantly related. The QTT framework effectively addresses these issues by enabling an exponential increase in discretisation points.
@@
Due to the intrinsic nonlinearity of the GPE, standard MPS methods, such as DMRG and TDVP, cannot be applied directly without modifications. We extend existing MPS algorithms to incorporate the nonlinearities of the GPE, allowing us to efficiently compute its ground states and dynamical evolution. To find the ground states, we consider two numerical approaches: ITE [Modugn2003EPJD...22..235M, Minguzzi2004], and variational methods. We investigate the efficiency and accuracy of the ground-state solutions of a cylindrically harmonically trapped BEC in two dimensions with and without rotation. Rotating a BEC results in the formation of quantum vortices, which organise themselves into a triangular lattice format, similar to the Abrikosov lattice in superconductors [Abrikosov1957]. For real-time dynamics, we consider the breathing-mode excitation, a BEC width oscillation with a trapping-frequency-dependent frequency for a sufficiently large BEC [pitaevskii2016bose, kimura2003breathing, muruganandam2009fortran], which can be generated by a sudden quench in the interaction strength. We show that the breathing-mode excitation can be accurately found in the real-time dynamics.
@@
We note that similar ideas have been explored independently in Ref. [Niedermeier2025, 1DQTTGPE, MPSGPE], which appeared concurrently with our work.
@@ 2 Gross-Pitaevskii equation
GPE is a nonlinear Schrödinger equation describing the weakly interacting BEC under mean-field approximation. In this work, we consider BEC in a rotating harmonic trap [kumar2019c]. The Hamiltonian is given by
@@
$\hat{H}_{\mathrm{int}}=g|\psi(\mathbf{r},t)|^{2},$ (3)
$\int d^{3}r\,\psi^{*}(\mathbf{r})\psi(\mathbf{r})=1\;$; (4)
where $\hat{H}^{1}$ is the one-body Hamiltonian including the kinetic energy, the harmonic trapping potential $\frac{1}{2}m\omega^{2}r^{2}$, and the rotational term $\Omega\hat{L}_{z}$. $\hat{H}_{\mathrm{int}}$ is the interaction with strength $g=4\pi\hbar^{2}a_{s}N/m$, where $a_{s}$ is the $s$-wave scattering length, $N$ is the total particle number and $m$ is the particle mass. Throughout this work, we consider the dimensionless Hamiltonian by adopting the characteristic scales of quantum harmonic oscillator units, $\sqrt{\hbar/m\omega}$, $1/\omega$ and $\hbar\omega$ for the length, time and energy respectively (This is equivalent to set $\hbar=1$, $m=1$, and $\omega=1$). The condensate wavefunction is normalized to one. (See more detail in Appendix A).
The evolution of the wavefunction is governed by the time-dependent GPE
$i\hbar\frac{\partial\psi(\mathbf{r},t)}{\partial t}=\hat{H}\psi(\mathbf{r},t).$ (5)
@@
$E^{1}[\psi]=\int d^{3}\mathbf{r}\,\psi^{*}(\mathbf{r})\hat{H}^{1}\psi(\mathbf{r})\;$; (8)
$E_{\mathrm{int}}[\psi]=\frac{g}{2}\int d^{3}\mathbf{r}|\psi(\mathbf{r})|^{4}.$ (9)
@@ 3 Methods
In this section, we introduce two methods for computing the ground state of the GPE, including the ITE and a gradient descent method. For real- or imaginary-time evolution, a so-called TDVP method is a powerful method and is widely used with the MPS/QTT framework [PhysRevLett.107.070601, PhysRevB.94.165116]. GD is effective for variationally optimizing the ground state. Here, we explain the necessary modifications to both methods to handle the nonlinear interaction term in the Hamiltonian.
In all our simulations, we use periodic boundary conditions. However, due to the harmonic trapping potential, the wavefunction decays away from the center and has a tiny amplitude (in the order of $10^{-9}$) at the boundary. Therefore, the boundary effect is negligible. In determining the system size, we use the Thomas-Fermi approximation (TFA) [RevModPhys.71.463, Cooper01112008, pitaevskii2016bose] to estimate the radius of the ground state, and choose a domain that is large enough to ensure a vanishing wavefunction amplitude at the boundary.
@@ 3.1 Imaginary-time evolution: TDVP
ITE is a common strategy for computing the ground state of the GPE [Modugn2003EPJD...22..235M, Minguzzi2004] and can be applied to the QTT wavefunction with minor modifications. Given the wavefunction (Hamiltonian) as a QTT (QTTO), the time evolution of a small time step $\delta\tau$ can be performed by using the so-called TDVP algorithm, which projects the evolution onto the tangent space of the QTT/MPS manifold, and ensures evolution remains representable as an QTT/MPS with controlled bond dimension (see Ref.[10.21468/SciPostPhysLectNotes.7] for an introductory lecture note.) Although the Hamiltonian in QTTO representation is typically “long-range” in the sense that it couples qubits over large distances, TDVP operates directly on the QTTO Hamiltonian, which makes it particularly well suited for QTT time evolutions. The Hamiltonian of the GPE contains the one-body term $\hat{H}^{1}$ and the interaction term $\hat{H}_{\mathrm{int}}$. The interaction term $g|\psi(\mathbf{r})|^{2}$ depends on the wavefunction $\psi(\mathbf{r})$, and therefore needs to be updated at each step using the evolved wavefunction from the last step. The time step $\delta\tau$ is set to be $0.5\delta x^{2}$ to satisfy the Courant-Friedrichs-Lewy condition and ensure the stability of the convergence.
An essential step in the TDVP algorithm is to compute the effective Hamiltonian, where the required tensor contractions are shown in Fig. 1(a). The computational cost is proportional to $dD^{3}D_{W}+d^{2}D^{2}D_{W}^{2}$, where $d=2$ is the dimension for each qubit, and $D$ and $D_{W}$ are the bond dimensions of the wavefunction and the Hamiltonian. We keep $\hat{H}^{1}$ and $\hat{H}_{\mathrm{int}}$ as two QTTOs and compute their effective Hamiltonians separately. For $\hat{H}^{1}$, the bond dimension is fixed and small, and the corresponding computational cost is relatively low. However, the QTTO for the interaction has a bond dimension $D^{2}$ for an exact calculation, where $D$ is the bond dimension of the wavefunction. Since $D$ controls the accuracy of the wavefunction and can be large in a challenging system, we need to compress $\psi^{2}(\mathbf{r})$ from a bond dimension $D^{2}$ to a smaller bond dimension $\chi$ to achieve efficient computations (see Fig. 1(b)). The computational cost then depends on the efficiency of this compression. In this work, we perform the compression using the variational method by maximizing the overlap $\langle\psi^{2}|\psi^{2}_{\chi}\rangle$ [SCHOLLWOCK201196], where $\psi^{2}_{\chi}(\mathbf{r})$ is the compressed $\psi^{2}(\mathbf{r})$ with the bond dimension $\chi$. In practice, $\psi^{2}(\mathbf{r})$ typically has a structure similar to that of $\psi(\mathbf{r})$, and therefore $\chi$ is typically on the same order as $D$. This yields an overall computational complexity of roughly $O(D^{4})$.
@@ 3.2 Variational method: gradient descent
Another powerful approach to the ground state is to use the variational method to minimize the energy $E[\psi]$. Although DMRG is not applicable due to the nonlinearity of the Hamiltonian, gradient descent remains a viable optimization method. Gradient descent has been used to optimize an MPS when the tensors are isometry tensors [10.21468/SciPostPhys.10.2.040, Zhu2017, ABRUDAN20091704, PhysRevB.107.085148]. In this work, we employ gradient descent within the MPS/QTT framework. The algorithm proceeds as follows. The QTT wavefunction is represented in canonical form, with one tensor at the orthogonality center–initially at the first site–and all tensors to the left (right) being left (right) orthogonal. The tensor at the orthogonality center is updated based on its energy gradient until it converges to a minimum. Once optimized, the orthogonality center is moved to the next site, and the process is repeated. This sweep continues back and forth until the global energy minimum is reached.
@@
Figure 2: Absolute errors in the gradient of the interaction energy (black curve) and in the slope along the descent direction (red curve) as functions of the number of tensors updated in $\psi_{\chi}^{2}$ after updating a tensor in $\psi$.